If z=8(cos5π/3+isin5π/3), then z6 is equal to


1
Expert's answer
2022-06-12T14:55:03-0400
"z^6=8^6\big(\cos(6(\dfrac{5\pi}{3}))+i\sin(6(\dfrac{5\pi}{3}))\big)""=262144\big(\cos(10\pi)+i\sin(10\pi)\big)=262144"